Aspen Construction made a basket purchase of three items. Item X was appraised at $ 36 comma 000 $36,000​; item Y was appraised at $ 60 comma 000 $60,000 and item Z was appraised at $ 63 comma 000 $63,000. The purchase price was $ 129 comma 000 $129,000. The amount at which item Y should be recorded​ is
Answer #1

Total appraised value = 36000+60000+63000 = 159000

Amount allocated to Item Y = 60000/159000*129000 = $48679